a. Suppose the natural rate of unemployment is 4 percent. What is the actual rate of unemployment if actual output is 2 percent below potential output? (Please express your answer in percent terms)

b. In Macroland, potential output equals $100 trillion and the natural rate of unemployment is 4 percent. If the actual unemployment rate is 3 percent, then real GDP equals:
Answer #1

As per Okun law, each percentage point of cyclical unemployment is associated with a 2 percent increase in the output gap

a)When output gap is 2% and the natural rate of unemployment is 4% then

Actual rate of unemployment is 5%

b) Potential output = 100 trillion

Natural unemployment = 4%

Actual unemployment = 3%

Cyclical unemployment = 4-3=1%

This means that the Real GDP is equal to 100+2 = $102 trillion